# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 3.3. Stability Condition
Suppose that $X$ is smooth and projective. We fix an ample line bundle $H$ on $X$ and define $H^{\prime}:=\pi_{X/k}^{*}H$, which is also an ample line bundle on $X^{\prime}$. Furthermore, $H^{p}\cong F_{X/k}^{*}H^{\prime}=F_{X}^{*}H$. Note that the stability condition of Higgs bundles on $X^{\prime}$ is given by the ample line bundle $H^{\prime}$ while the stability condition of flat bundles on $X$ is given by $H$. In this subsection, we improve the equivalence given in Corollary 3.3 by adding the information of stability conditions:
@@
There is an equivalence of categories between ${\rm HIG}^{(s)s}_{0}(X^{\prime}/k,G)$ and ${\rm MIC}^{(s)s}_{0}(X/k,G)$.
@@
Let $E$ be a $G$-bundle on $X^{\prime}$, and denote by $(V,\nabla_{can})$ the corresponding flat $G$-bundle on $X$ with trivial $p$-curvature by Corollay 3.3, where $V=F^{*}_{1}E$. Let $P$ be a parabolic subgroup of $G$. To prove the equivalence of stability conditions, it is enough to prove the following two statements:
@@
    there is a one-to-one correspondence between reductions of structure group $\sigma:X^{\prime}\rightarrow E/P$ and $\nabla_{can}$-compatible reductions of structure group $\varsigma:X\rightarrow V/P$;
@@
    $p\deg_{H}L(\varsigma,\chi)=\deg_{H^{\prime}}L(\sigma,\chi),$
@@
For the first statement, given a reduction of structure group $\sigma:X^{\prime}\rightarrow E/P$, we obtain a reduction $\varsigma:X\rightarrow V/P$ by pullback under the relative Frobenius morphism $F_{1}$. In other words, the morphism $E_{\sigma}\rightarrow E$ induces a morphism $(V_{\varsigma},\nabla_{can,\varsigma})\rightarrow(V,\nabla_{can})$ of flat $G$-bundles by the classical Cartier descent (Theorem 3.2), where $V_{\varsigma}=F^{*}_{1}E_{\sigma}$ and $\nabla_{can,\varsigma}$ is the canonical connection on $V_{\varsigma}$ given by Cartier descent. Therefore, the diagram
${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes\Omega_{X}}$${\mathcal{O}_{V_{\varsigma}}}$${\mathcal{O}_{V_{\varsigma}}\otimes\Omega_{X}}$$\scriptstyle{\nabla_{can}}$$\scriptstyle{\nabla_{can,\varsigma}}$
commutes. This shows that $\varsigma$ is $\nabla_{can}$-compatible. By the classical Cartier descent, it is easy to see that this is a one-to-one correspondence. The second statement is clear because $H^{p}\cong F^{*}_{1}H^{\prime}$ and $V_{\varsigma}=F^{*}_{1}E_{\sigma}$. ∎
Now we consider the case of nilpotency of exponent $\leq p-1$.
@@
Let $(E,\theta)\in{\rm HIG}_{p-1}(X^{\prime}/k,G)$ be a $G$-Higgs bundle, and denote by $(V,\nabla)\in{\rm MIC}_{p-1}(K/k,G)$ the corresponding flat $G$-bundle under the equivalence given by Theorem 3.1. To prove the theorem, it is enough to show that $\theta$-compatible reductions $\sigma$ of $E$ are in one-to-one correspondence with $\nabla$-compatible reductions $\varsigma$ of $V$.
@@ 4. Logarithmic G-Higgs bundles and Flat G-bundles on Root Stacks
In this section, we are mostly interested in a special classes of smooth Deligne–Mumford stacks $\mathscr{X}$ considered and discussed in [MO05, Cad07, Sim11], which are called root stacks in this paper (Theorem 4.3). We define logarithmic nilpotent $G$-Higgs bundles and logarithmic nilpotent flat $G$-bundles on $\mathscr{X}$ of exponent $\leq p-1$ and prove an equivalence of categories between ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ under a $W_{2}$-lifting condition (Theorem 4.12). Then, we follow Borne and Nironi’s approach to give the stability condition [Bor07, Nir09], and prove the equivalence of categories under stability conditions (Theorem 4.16).
@@ 4.1. G-Higgs Bundles and Flat G-bundles on DM stacks
Recall that a Deligne–Mumford stack over $k$ is an algebraic stack such that there exists a surjective étale morphism $Y\rightarrow\mathscr{X}$, where $Y$ is a scheme. It is well-known that $\mathscr{X}\cong[Y/R]$, where $R=Y\times_{\mathscr{X}}Y$ is an algebraic space and we have two étale morphisms
@@
for $i=1,2$. For convenience, we suppose that $R$ is a scheme. In fact, we can apply all arguments in this section to algebraic spaces first, and then give the proof for Deligne–Mumford stacks. Let $F_{\mathscr{X}/k}:\mathscr{X}\rightarrow\mathscr{X}^{\prime}$ be the relative Frobenius morphism, for which we still use the notation $F_{1}:=F_{\mathscr{X}/k}$ for simplicity, and we suppose that $\mathscr{X}$ is $W_{2}(k)$-liftable. Here is an equivalent description under the isomorphism $\mathscr{X}\cong[Y/R]$. The relative Frobenius morphism $F_{\mathscr{X}/k}$ is induced from those on $Y$ and $R$, and $\mathscr{X}$ is $W_{2}(k)$-liftable is equivalent to say that $Y$ and $R$ are $W_{2}(k)$-liftable and the following diagrams commutes
@@
The stack $\mathscr{X}$ is called tame if the induced functor
@@
Suppose that $\mathscr{X}$ is tame, smooth and $W_{2}(k)$-liftable with coarse moduli space $\pi:\mathscr{X}\rightarrow X$, where $X$ is a smooth variety over $k$. Denote by $\Omega_{\mathscr{X}/k}$ the relative cotangent sheaf, and if there is no confusion, we use the notation $\Omega_{\mathscr{X}}$ for simplicity. Suppose that $\mathscr{X}\cong[Y/R]$, and then a $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ can be regarded as a pair $(E,\tau)$, where $E$ is a $G$-bundle on $Y$ and $\tau:p_{1}^{*}E\xrightarrow{\cong}p_{2}^{*}E$ is an isomorphism. In the same way, a $G$-Higgs bundle $(\mathscr{E},\theta)$ on $\mathscr{X}$ is regarded as a triple $(E,\theta,\tau)$ on $Y$, where $(E,\theta)$ is a Higgs bundle on $Y$ and $\tau:p_{1}^{*}(E,\theta)\rightarrow p_{2}^{*}(E,\theta)$ is an isomorphism, and a similar description can be given for flat $G$-bundles $(\mathscr{V},\nabla)$ on $\mathscr{X}$. Therefore, the nilpotency for $G$-Higgs bundles and flat $G$-bundles on $\mathscr{X}$ is defined as we did on varieties in §2.2. Define the follow categories:
  • ${\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)$ is the category of $G$-bundles on $\mathscr{X}^{\prime}$,
@@
  • ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}/k,G)$ is the category of nilpotent $G$-Higgs bundles on $\mathscr{X}^{\prime}$ of exponent $\leq p-1$,
@@
Moreover, in the rest of the paper, we only give the definition of categories of objects with nilpotent exponent $\leq p-1$ and omit the deifnition for objects with nilpotent exponent $\leq 0$ for simplicity.
@@
where $F_{1}$ is the relative Frobenius morphism. Given a $G$-bundle $\mathscr{E}$ on $\mathscr{X}^{\prime}$, it can be regarded as a pair $(E,\tau)$, where $E$ is a $G$-bundle on $Y^{\prime}$ and $\tau:p_{1}^{*}E\cong p_{2}^{*}E$ is an isomorphism of $G$-bundles on $R^{\prime}$. Since $F_{1}\circ p_{i}=p_{i}\circ F_{1}$, we have $F_{1}^{*}p_{i}^{*}E\cong p_{i}^{*}F_{1}^{*}E$. By Corollary 3.3, the $G$-bundle $E$ corresponds to a flat $G$-bundle $(V,\nabla_{can})$ on $Y$, where $V=F^{*}_{1}E$. On the other hand, the $G$-bundle $p_{i}^{*}E$ on $R^{\prime}$ corresponds to a flat $G$-bundle on $R$, which is isomorphic to $p_{i}^{*}(V,\nabla_{can})$ by the isomorphism $F_{1}^{*}p_{i}^{*}E\cong p_{i}^{*}F_{1}^{*}E$ and uniqueness of the canonical connection. Thus, we obtain an isomorphism $p_{1}^{*}(V,\nabla_{can})\cong p_{2}^{*}(V,\nabla_{can})$ induced from $\tau:p_{1}^{*}E\rightarrow p_{2}^{*}E$, and denote it by $F_{1}^{*}\tau$. The data $((V,\nabla_{can}),F^{*}\tau)$ gives a flat $G$-bundle $(\mathscr{V},\nabla_{can})$ with trivial $p$-curvature on $\mathscr{X}$. The other direction can be proved similarly. ∎
@@
The proof is similar to that of Proposition 4.1, and we still work on $\mathscr{X}\cong[Y/R]$. By Theorem 3.1, we have the following equivalences of categories:
  • ${\rm HIG}_{p-1}(Y^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(Y/k,G)$;
  • ${\rm HIG}_{p-1}(R^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(R/k,G)$.
Given a $G$-Higgs bundle $(\mathscr{E},\theta)$ on $\mathscr{X}^{\prime}$, we regard it as a $G$-Higgs bundle $(E,\theta)$ on $Y^{\prime}$ together with an isomorphism $\tau$. The first equivalence of categories gives a flat $G$-bundle $(V,\nabla)\in{\rm MIC}_{p-1}(Y/k,G)$, and then we obtain flat $G$-bundles $p_{i}^{*}(V,\nabla)\in{\rm MIC}_{p}(R/k,G)$ for $i=1,2$. Via étale morphisms $p_{i}$ for $i=1,2$, we obtain $G$-Higgs bundles on $R^{\prime}$, which are isomorphic to $p_{i}^{*}(E,\theta)$ by the commutative diagram in Proposition 4.1. Therefore, the isomorphism $\tau$ induces an isomorphism $p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla)$, and denote it by $F^{*}\tau$. Thus, the data $((V,\nabla),F^{*}\tau)$ gives a flat $G$-bundle in ${\rm MIC}_{p-1}(\mathscr{X}/k,G)$. The proof of the other direction is given in the same way. ∎

### v2

@@ 3.3. Stability condition
Suppose that $X$ is smooth and projective. We fix an ample line bundle $H$ on $X$. Define $H^{\prime}:=\pi_{X/k}^{*}H$, which is also an ample line bundle on $X^{\prime}$. Furthermore, $H^{p}\cong F_{X/k}^{*}H^{\prime}\cong F_{X}^{*}H$. We remind the reader that the stability condition of $G$-Higgs bundles on $X^{\prime}$ is given by the ample line bundle $H^{\prime}$, while the stability condition of flat $G$-bundles on $X$ is given by $H$.
We first improve the equivalence given in Corollary 3.3 by adding the information of stability conditions:
@@
The category ${\rm HIG}^{(s)s}_{0}(X^{\prime}/k,G)$ is equivalent to ${\rm MIC}^{(s)s}_{0}(X/k,G)$.
@@
Let $E$ be a $G$-bundle on $X^{\prime}$, and denote by $(V,\nabla_{can})$ the corresponding flat $G$-bundle on $X$ with trivial $p$-curvature by Corollary 3.3, where $V=F^{*}_{1}E$. Let $P$ be a parabolic subgroup of $G$. To prove the equivalence of stability conditions, it is enough to prove the following two statements:
@@
    there is a one-to-one correspondence between reductions of structure group $\sigma:(X^{\rm big})^{\prime}\rightarrow(E|_{(X^{\rm big})^{\prime}})/P$ and $\nabla_{can}$-compatible reductions of structure group $\varsigma:X^{\rm big}\rightarrow(V|_{X^{\rm big}})/P$;
@@
    $\deg_{H}L(\varsigma,\chi)=p\deg_{H^{\prime}}L(\sigma,\chi),$
@@
For the first statement, although the stability condition (Definition 2.4 and 2.5) is given for reductions of structure group over $X^{\rm big}$, it is enough to work on $X$. given a reduction of structure group $\sigma:X^{\prime}\rightarrow E/P$, we obtain a reduction $\varsigma:X\rightarrow V/P$ via the relative Frobenius morphism $F_{1}$. Equivalently, the morphism $E_{\sigma}\rightarrow E$ induces a morphism $(V_{\varsigma},\nabla_{can,\varsigma})\rightarrow(V,\nabla_{can})$ of flat $G$-bundles by the classical Cartier transform (Theorem 3.2), where $V_{\varsigma}=F^{*}_{1}E_{\sigma}$. Therefore, $\varsigma$ is $\nabla_{can}$-compatible. By Theorem 3.2, the correspondence between reduction of structure groups $\sigma$ and $\varsigma$ is a one-to-one correspondence. The second statement is clear with respect to the correspondence given in the first statement. ∎
Now we consider the case of nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles of exponent $\leq p-1$.
@@
Let $(E,\theta)\in{\rm HIG}_{p-1}(X^{\prime}/k,G)$ be a $G$-Higgs bundle, and denote by $(V,\nabla)\in{\rm MIC}_{p-1}(K/k,G)$ the corresponding flat $G$-bundle under the equivalence given by Theorem 3.1. With the same approach as Proposition 3.7, it is enough to show that $\theta$-compatible reductions $\sigma:X^{\prime}\rightarrow E/P$ are in one-to-one correspondence with $\nabla$-compatible reductions $\varsigma:X\rightarrow V/P$.
@@ 4. Logarithmic G-Higgs bundles and Flat G-bundles on Root Stacks
In this section, we are mostly interested in a special classes of smooth Deligne–Mumford stacks $\mathscr{X}$ considered and discussed in [MO05, Cad07, Sim11], which are called root stacks in this paper (Theorem 4.3). We prove an equivalence of categories between logarithmic nilpotent $G$-Higgs bundles and logarithmic nilpotent flat $G$-bundles of exponent $\leq p-1$ on root stacks (Theorem 4.12), which also preserves stability conditions (Theorem 4.15).
@@ 4.1. G-Higgs Bundles and Flat G-bundles on DM stacks
Recall that a Deligne–Mumford stack over $k$ is an algebraic stack such that there exists a surjective étale morphism $Y\rightarrow\mathscr{X}$, where $Y$ is a scheme. It is well-known that $\mathscr{X}\cong[Y/R]$, where $R=Y\times_{\mathscr{X}}Y$ is an algebraic space together with two étale morphisms
@@
for $i=1,2$. For convenience, we suppose that $R$ is a scheme. In fact, we can apply all arguments in this section to algebraic spaces first, and then move to Deligne–Mumford stacks. Let $F_{1}:=F_{\mathscr{X}/k}:\mathscr{X}\rightarrow\mathscr{X}^{\prime}$ be the relative Frobenius morphism and we suppose that $\mathscr{X}$ is $W_{2}(k)$-liftable. Here is an equivalent description under the isomorphism $\mathscr{X}\cong[Y/R]$. The relative Frobenius morphism $F_{\mathscr{X}/k}$ is induced from those on $Y$ and $R$, and the condition that $\mathscr{X}$ is $W_{2}(k)$-liftable is equivalent to say that $Y$ and $R$ are $W_{2}(k)$-liftable and the following diagram commutes
@@
where $\widetilde{p}_{i}$ is an étale morphism for $i=1,2$. The stack $\mathscr{X}$ is called tame if the induced functor
@@
Suppose that $\mathscr{X}$ is tame, smooth and $W_{2}(k)$-liftable with coarse moduli space $\pi:\mathscr{X}\rightarrow X$, where $X$ is a smooth variety over $k$. Denote by $\Omega_{\mathscr{X}}:=\Omega_{\mathscr{X}/k}$ the relative cotangent sheaf. Given $\mathscr{X}\cong[Y/R]$, a $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ can be regarded as a pair $(E,\tau)$, where $E$ is a $G$-bundle on $Y$ and $\tau:p_{1}^{*}E\xrightarrow{\cong}p_{2}^{*}E$ is an isomorphism satisfying the cocycle condition on $Y\times_{\mathscr{X}}Y\times_{\mathscr{X}}Y$. In the same way, a $\lambda$-connection on $\mathscr{X}$ is regarded as a triple $(E,\nabla,\tau)$ on $Y$, where $(E,\nabla)$ is a $\lambda$-connection on $Y$ and $\tau:p_{1}^{*}(E,\nabla)\rightarrow p_{2}^{*}(E,\nabla)$ is an isomorphism satisfying the cocycle condition. The nilpotency for $G$-Higgs bundles and flat $G$-bundles on $\mathscr{X}\cong[Y/R]$ is defined as the nilpotency of the corresponding $G$-Higgs bundles and flat $G$-bundles on $Y$. Note that this definition does not depend on the choice of the surjective étale morphism $Y\rightarrow\mathscr{X}$. We define the following categories:
  • ${\rm HIG}_{0}(\mathscr{X}/k,G)$ is the category of $G$-bundles on $\mathscr{X}$,
@@
  • ${\rm HIG}_{p-1}(\mathscr{X}/k,G)$ is the category of nilpotent $G$-Higgs bundles on $\mathscr{X}$ of exponent $\leq p-1$,
@@
where $F_{R,1}$ and $F_{Y,1}$ are the relative Frobenius morphisms. Given a $G$-bundle $\mathscr{E}$ on $\mathscr{X}^{\prime}$, it can be regarded as a pair $(E,\tau^{\prime})$, where $E$ is a $G$-bundle on $Y^{\prime}$ and $\tau^{\prime}:(p^{\prime}_{1})^{*}E\cong(p^{\prime}_{2})^{*}E$ is an isomorphism of $G$-bundles on $R^{\prime}$ satisfying the cocycle condition. Since $F_{Y,1}\circ p=p^{\prime}\circ F_{R,1}$, where $p=p_{1},p_{2}$, we have $F_{R,1}^{*}p^{\prime*}E\cong p^{*}F_{Y,1}^{*}E$. By Corollary 3.3, the $G$-bundle $E$ on $Y^{\prime}$ corresponds to a flat $G$-bundle $(V,\nabla_{can})$ on $Y$, where $V=F^{*}_{Y,1}E$. On the other hand, the $G$-bundle $p^{\prime*}E$ on $R^{\prime}$ corresponds to a flat $G$-bundle on $R$. This flat $G$-bundle is isomorphic to $p^{*}(V,\nabla_{can})$ by the isomorphism $F_{R,1}^{*}p^{\prime*}E\cong p^{*}F_{Y,1}^{*}E$ and the uniqueness of the canonical connection. Thus, we obtain an isomorphism $\tau:p_{1}^{*}(V,\nabla_{can})\cong p_{2}^{*}(V,\nabla_{can})$ induced from $\tau^{\prime}$. The data $((V,\nabla_{can}),\tau)$ gives a flat $G$-bundle $(\mathscr{V},\nabla_{can})$ with trivial $p$-curvature on $\mathscr{X}$. The other direction can be proved in the same way. ∎
@@
The proof of this theorem is a stacky version of that of Theorem 3.1. We give a proof under the isomorphism $\mathscr{X}\cong[Y/R]$. Before we move to construct inverse Cartier and Cartier transform, we first enrich Lemma 3.4 under étale morphisms. We have the following diagram
[…]
where
  • $p$ and $\widetilde{p}$ are regarded as étale morphisms $p_{i}$ and $\widetilde{p}_{i}$ respectively, and $p^{\prime}$ and $\widetilde{p}^{\prime}$ are the induced ones;
  • $F_{Y,1}$ and $F_{R,1}$ are relative Frobenius morphisms, and $F_{Y,2}$ and $F_{R,2}$ are Frobenius liftings, of which mod $p$ are $F_{Y,1}$ and $F_{R,1}$ respectively.
We fix an open cover $\{U_{Y,i}\}_{i\in I}$ of $Y$ and denote by $\{U_{R,i}\}_{i\in I}$ the corresponding open cover of $R$ under the étale morphism $p$. By Lemma 3.4, we have
$\zeta_{\bullet,i}:=\frac{F^{*}_{\bullet,2}}{[p]}:F^{*}_{\bullet,1}\Omega_{U^{\prime}_{\bullet,i}}\rightarrow\Omega_{U_{\bullet,i}}$
$h_{\bullet,ij}:F^{*}_{\bullet,1}\Omega_{U^{\prime}_{\bullet,ij}}\rightarrow\Omega_{U_{\bullet,ij}}$
such that
$\zeta_{\bullet,i}-\zeta_{\bullet,i}=dh_{\bullet,ij}$
$h_{\bullet,ij}+h_{\bullet,jk}=h_{\bullet,ik},$
where $\bullet=Y\text{ or }R$. Since $F_{Y,2}\circ\widetilde{p}=\widetilde{p}^{\prime}\circ F_{R,2}$ and $F_{Y,1}\circ p=p^{\prime}\circ F_{R,1}$, the morphisms
$\frac{\widetilde{p}^{*}F^{*}_{Y,2}}{[p]}:p^{*}F^{*}_{Y,1}\Omega_{U^{\prime}_{Y,i}}\rightarrow p^{*}\Omega_{U_{Y,i}}\cong\Omega_{U_{R,i}}$
and
$\frac{F^{*}_{R,2}\widetilde{p}^{\prime*}}{[p]}:F^{*}_{R,1}p^{\prime*}\Omega_{U^{\prime}_{Y,i}}\rightarrow\Omega_{U_{R,i}}$
are the same, i.e.
(4.1) $\frac{\widetilde{p}^{*}F^{*}_{Y,2}}{[p]}=\frac{F^{*}_{R,2}\widetilde{p}^{\prime*}}{[p]}.$
Now given a $G$-Higgs bundle on $\mathscr{X}^{\prime}$, we regard it as triple a $(E,\theta,\tau^{\prime})$, where $(E,\theta)$ is a $G$-Higgs bundle on $Y^{\prime}$ and $\tau^{\prime}:(p^{\prime}_{1})^{*}(E,\theta)\cong(p^{\prime}_{2})^{*}(E,\theta)$ is an isomorphism satisfying the cocycle condition. The inverse Cartier transform in §3.1 gives a flat $G$-bundle $(V,\nabla):=C^{-1}_{\rm exp}(E,\theta)$ on $Y$. We will construct an isomorphism $\tau:p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla)$ satisfying the cocycle condition and thus obtain a flat $G$-bundle on $\mathscr{X}$. Since $p^{\prime}_{i}:R^{\prime}\rightarrow Y^{\prime}$ is étale, $(p^{\prime}_{i})^{*}(E,\theta)$ is a $G$-Higgs bundle on $R^{\prime}$. Applying the inverse Cartier transform $C^{-1}_{\rm exp}$ on $R^{\prime}$ again, we obtain flat $G$-bundles $C^{-1}_{\rm exp}((p^{\prime}_{i})^{*}(E,\theta))$ on $R$ for $i=1,2$ together with an isomorphism
$C^{-1}_{\rm exp}(\tau^{\prime}):C^{-1}_{\rm exp}((p^{\prime}_{1})^{*}(E,\theta))\cong C^{-1}_{\rm exp}((p^{\prime}_{2})^{*}(E,\theta)).$
We claim that
$C^{-1}_{\rm exp}((p^{\prime})^{*}(E,\theta))\cong p^{*}(V,\nabla)=p^{*}(C^{-1}_{\rm exp}(E,\theta)),$
where $p=p_{1},p_{2}$, and thus we obtain the desired isomorphism $\tau:=C^{-1}_{\rm exp}(\tau^{\prime})$ satisfying the cocycle condition.
Although the claim has been proved in [She24], we include the proof here because the construction of the Cartier transform is exactly given in the same way. Let $(E_{i},\theta_{i})$ be the restriction of $(E,\theta)$ on $U^{\prime}_{Y,i}$. The flat $G$-bundle $p^{*}(C^{-1}_{\rm exp}(E,\theta))$ is obtained by gluing the local data
$(p^{*}F^{*}_{Y,1}E_{i},\ \nabla_{can,i}+p^{*}\zeta_{Y,i}(F^{*}_{Y,1}\theta_{i}))_{i\in I}$
via
$\overline{\rm exp}(p^{*}h_{Y,ij}(F^{*}_{Y,1}\theta_{ij})).$
Similarly, $C^{-1}_{\rm exp}((p^{\prime})^{*}(E,\theta))$ is obtained by gluing the local data
$(F^{*}_{R,1}p^{\prime*}E_{i},\ \nabla_{can,i}+\zeta_{R,i}(F^{*}_{R,1}p^{\prime*}\theta_{i}))_{i\in I}$
via
$\overline{\rm exp}(h_{R,ij}(F^{*}_{R,1}p^{\prime*}\theta_{ij})).$
Clearly,
$p^{*}F^{*}_{Y,1}E_{i}=F^{*}_{R,1}p^{\prime*}E_{i}.$
Then, by (4.1), we have
$p^{*}\zeta_{Y,i}(F^{*}_{Y,1}\theta_{i})=\zeta_{R,i}(F^{*}_{R,1}p^{\prime*}\theta_{i})$
and
$p^{*}h_{Y,ij}(F^{*}_{Y,1}\theta_{ij})=h_{R,ij}(F^{*}_{R,1}p^{\prime*}\theta_{ij}).$
Therefore,
$(p^{*}F^{*}_{Y,1}E_{i},\ \nabla_{can,i}+p^{*}\zeta_{Y,i}(F^{*}_{Y,1}\theta_{i}))=(F^{*}_{R,1}p^{\prime*}E_{i},\ \nabla_{can,i}+\zeta_{R,i}(F^{*}_{R,1}p^{\prime*}\theta_{i}))$
and
$\overline{\rm exp}(p^{*}h_{Y,ij}(F^{*}_{Y,1}\theta_{ij}))=\overline{\rm exp}(h_{R,ij}(F^{*}_{R,1}p^{\prime*}\theta_{ij})).$
This finishes the proof for the claim.
The construction of the Cartier transform is exactly the same and we leave it for the reader. ∎